Cheetahs, the fastest of the great cats can reach 20.1 m/s in 2.0s starting from rest. Assuming that they have constant acceleration throughout that time, find (a) their acceleration and (b) the distance they travel during that time
Answer #1

Given : vf = 20.1 m/s , v0 = 0 m/s , t = 2 s

Solution :

(a) Acceleration

Using the equation of motion :

vf = vi + at

20.1 m/s = 0 + a(2s)

= 10.05 m/s2

Answer : 10.05 m/s2

(b) Distance travelled

Using equation of motion :

d = v0t + (1/2)at2

= (0)(2s) + (1/2)(10.05 m/s2)(2s)2

= 20.1 m

Answer : 20.1 m